OpenStudy (anonymous):

Simplify: -7[-44 - (-97 - 37)] =

OpenStudy (anonymous):

-630

OpenStudy (anonymous):

-7[-44 - (-97 - 37)] = -7[-44 - (-134)] = -7[-44 + 134] = -7[90] = -630

OpenStudy (anonymous):

-7[-44-(-134)] -7[-44+134] -7[90] -630